OpenStudy (anonymous):

Evaluate cos(tan^-1 0)

zepdrix (zepdrix):

\[\large \cos(\arctan 0)\] Hmm ok so let's first look at the inner part, \(\arctan 0\). Since it's an inverse trig function, the solution is actually an angle. We can write it like this.\[\large \arctan 0 = \theta\] So our problem is actually,\[\large \cos(\arctan 0) \qquad \rightarrow \qquad \cos(\theta)\] So first of all what is the arctan of zero? If you're not sure, you could rewrite as tangent undoing the inverse. \[\large \arctan 0 =\theta \qquad \rightarrow \qquad \tan \theta= 0\]This occurs when theta is zero. So our problem is now,\[\large \cos(\arctan 0) \qquad \rightarrow \qquad \cos(0)\]Remember your special angle for this one? :)

OpenStudy (zehanz):

So:\[\cos (\tan^{-1} (0))\] You probably know: tan 0 =0. That gives you tan^-1(0) as well.

OpenStudy (anonymous):

`right

OpenStudy (zehanz):

So tan^-1(0)=0 and you are to find cos(0)=...?

OpenStudy (anonymous):

1

OpenStudy (zehanz):

Done!